how do i find the horizontal distance on this problem:
data:
vi 8m/s
dx= ?
a=om/s
t=3.061

i tried dx=8m/s(3.061) and obtained 24.488
is dx=24.488?

Answer by Alan3354(69443)

how do i find the horizontal distance on this problem:
data:
vi 8m/s
dx= ?
a=om/s
t=3.061

i tried dx=8m/s(3.061) and obtained 24.488
is dx=24.488?
--------------
If Vi is the horizontal component of the velocity, then it's
s = vt
s = 8*3.061 = 24.488 meters
----------
You use Vi, then dx. Both are commonly used for horizontal, but not in the same problem.